OpenStudy (anonymous):

P and Q are two buoys, Q being 250 metres due east of P.If a ship 250metres N.62 degree E. from P, calculate her distance from Q.

OpenStudy (anonymous):

wait is the distance form p to the ship 250?

OpenStudy (anonymous):

*from p to the ship

OpenStudy (anonymous):

then you could use the law of cosines look at my attachment

OpenStudy (anonymous):

in this case B and C are both 250m and a is 62 degrees so just plug those into the law of cosines to solve for A^2 and then take the square root to solve for A

OpenStudy (anonymous):

 its this right? yah then u can use the law of cosines

OpenStudy (anonymous):

so using the law of cosines: \[A^2=250^2+250^2-2*250*250*\cos(62)\]\and then take the square root of that to solve for A

OpenStudy (anonymous):

law of cosines:\[A^2=B^2+C^2-2BCcos(a)\]

OpenStudy (anonymous):

so the answer is 257.5 meters